A bangka (boat) starts from rest. After 10 seconds, the bangka has a speed of 40 km/h.

(a) Compute the bangka’s final velocity in units of m/s.

(b) Compute the bangka’s average acceleration.


1
Expert's answer
2021-11-15T10:33:09-0500

"v_0=0 \frac{m}{s} \space\space\space\space t = 10s \space\space\space\space v=40\frac{km}{h}"

a) the final velocity is "v=" "40\frac{km}{h} = 40 * \frac{5}{18}=\frac{100}{9}=11\frac{1}{9}\frac{m}{s}"

b) average acceleration "a = \large\frac{v-v_o}{t}" "=\frac{100}{9}*\frac{1}{10}=\frac{10}{9}\frac{m}{s}=1\frac{1}{9}\frac{m}{s^2}"